in an X ray machine, an electron is subjected to force of 10-23N.In how much time will the electron cover a distance of 0.1m? mass of electron=10-30kg.

Grade:9

1 Answers

Eshan
5 years ago
Dear student,

a=\dfrac{F}{m}
s=\dfrac{1}{2}at^2\implies t=\sqrt{\dfrac{2s}{a}}=\sqrt{\dfrac{2sm}{F}}

=0.141ms